# : Context Management, Validation, and Transaction Guarantees for Multi-Agent LLM Planning

### v1

@@ 5.4. Observations
Our experiments highlight key limitations in standalone LLMs for complex planning. The most critical is attention narrowing, where models lose sight of global constraints, miscalculate travel times, and fail to audit their plans (self-validation failure).
Common-sense gaps emerge in task handling, with models arbitrarily inserting, omitting, or misjudging time for essential steps (e.g., luggage retrieval, walking). While sometimes negligible, such errors can be critical.
LLMs also struggle with state management, often βrewritingβ past events rather than adapting, leading to inconsistent spatial-temporal reasoning. Some models lost track of agent locations, producing infeasible plans.
Our results further show varied performance across LLMs. GPT-o1 maintained partial historical awareness to some extent, though it still exhibited lapses. DeepSeekΒ R1, ClaudeΒ 3.7, and GPT-4o all showed significant stateless and context-retention issues in reactive planning scenarios.
By contrast, the framework mitigates these issues, as discussed in our conclusion.
@@ 6. Conclusion
This paper introduced , a structured multi-agent framework designed to address fundamental limitations in monolithic LLM-based planning and reasoning systems. Through systematic experimentation using the REALM benchmark, we identified and remediated four critical shortcomings in current approaches: inadequate self-validation, context narrowing, lack of transaction properties, and insufficient inter-agent coordination. Our findings demonstrate that even state-of-the-art LLMs such as Claude 3.7 and GPT-o1, despite their impressive capabilities, struggle to maintain global constraint awareness during complex planning tasks. They exhibit a tendency to fixate on recent adjustments while progressively disregarding earlier constraints, leading to inconsistent and infeasible plans. This attention narrowing effect is further exacerbated during reactive planning scenarios, where models attempt to rewrite already-executed actions rather than adapting from the current state. The framework addresses these limitations through several key innovations:
1.
Structured validation protocols: Enforce constraint checks at critical decision points.
2.
Context management mechanisms: Checkpoint historical state transitions and constraints to prevent catastrophic context loss.
3.
Transaction state preservation: Maintain immutable historical actions while enabling adaptive future planning.
4.
Specialized agent distribution: Assign focused responsibilities to prevent context overload and manage inter-agent dependency resolution.
These components work together to enable robust planning and reasoning in diverse scenarios, from complex travel logistics to time-sensitive activities with multiple interdependencies. By distributing the cognitive load among specialized agents and imposing rigorous validation, maintains consistency and adaptability even under significant disruptions.
Our findings suggest that the future of LLM-based systems lies not in scaling up monolithic models but in orchestrating smaller specialized agents through well-designed coordination mechanisms. This approach boosts performance on complex tasks while improving predictability, reliability, and transparency, all vital for mission-critical applications.
Future research will focus on mitigating context narrowing within autoregressors, developing formal verification methods for multi-agent systems, integrating with external storage for managing context (Huang etΒ al., 2020), and extending to address scientific reasoning, creative collaboration, and complex decision making under uncertainty.

### v2

@@ 5.4. Observations
Our experiments across multiple LLMs (GPT-o1, DeepSeekΒ R1, ClaudeΒ 3.7, GPT-4o) highlight consistent limitations in complex planning scenarios. While GPT-o1 showed partial historical awareness, all models exhibited attention narrowing, self-validation failure, and inconsistent spatial-temporal reasoning.
TableΒ 9 summarizes βs context management and compensation mechanisms directly address these limitations of LLMs.
| Capability | Standard LLMs | |
| Maintains historical actions | Partial/None | Full |
| Partial journey compensation | Rarely | Always |
| Constraint consistency checking | Ad-hoc | Systematic |
| Handles attention narrowing | Vulnerable | Resistant |
| Physical-temporal consistency | Inconsistent | Guaranteed |
@@ 6. Conclusion
We introduced , a structured transactional multi-agent framework specifically designed to overcome fundamental weaknesses of monolithic LLM-based planning systems. Through systematic experimentation using the REALM benchmark, we identified and addressed four critical shortcomings of existing approaches: inadequate self-validation, context narrowing, absence of transaction properties, and insufficient inter-agent coordination.
Our results demonstrate that even state-of-the-art LLMs, such as Claude 3.7 and GPT-o1, despite strong reasoning capabilities, often fixate on recent context while neglecting critical earlier constraints, leading to inconsistent and infeasible plans. This issue is particularly pronounced in reactive planning scenarios, where models frequently attempt to retroactively rewrite past actions rather than adapting from the current state.
The framework addresses these limitations through four key innovations.
Structured validation protocols: Independent constraint checks to overcome inadequate self-validation.
Context management mechanisms: Strategic checkpointing to mitigate context narrowing in long sequences.
Transactional state preservation: Immutable historical records and compensatory mechanisms to address missing transaction properties.
Specialized agent distribution: Explicit role specialization and dependency tracking to resolve insufficient inter-agent coordination.
Collectively, these innovations enable robust and coherent planning in diverse real-world scenarios, from complex travel logistics to dynamic, time-sensitive tasks. By delegating tasks among specialized agents and enforcing rigorous transactional validation, significantly improves consistency, reliability, transparency, and adaptabilityβqualities essential for mission-critical applications.
Future research will address intrinsic context narrowing in autoregressive models, develop formal verification methods for transactional multi-agent systems, integrate advanced external context-management tools (Huang etΒ al., 2020), and extend to complex domains such as scientific reasoning, creative collaboration, and decision-making under uncertainty.
@@ Acknowledgment
Supplementary Materials
Benchmark: Our established benchmark (Geng and Chang, 2025) for evaluating multi-agent systems is located at
https://github.com/genglongling/REALM-Bench. -
Experiment Source Code: The source code of the for solving the four benchmark problems is located at
https://github.com/genglongling/SagaLLM.